OpenStudy (anonymous):

Who here is really good at math?????? ONLY TEN MINUTES LEFT, PLEASE HELP!!!!!!!!!!!!!! come to this question if you are and see if you can hopefully help me are 18 + 45 = ___ ´ (2 + 5) 2 3 7 9

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

questions

OpenStudy (anonymous):

does anyone here have free time like, an hour or 45 mins?

OpenStudy (anonymous):

i can do one.

OpenStudy (anonymous):

if you can leave

OpenStudy (anonymous):

cant

OpenStudy (anonymous):

i have 10 questions

OpenStudy (anonymous):

and nobody's going to help you with ten, or give you all the answers.

OpenStudy (anonymous):

Let's see :)

OpenStudy (anonymous):

@marihelenh helped me with 25 some people are just that nice, but they logged off

OpenStudy (anonymous):

meh i could do 10 but WE NEED TO SEE THE QUESTIONS

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

ok once i start i cant back out, here we go

OpenStudy (anonymous):

What value can be written on the blank line to make the expressions equivalent? 48 + 54 = ___ ´ (8 + 9)

OpenStudy (anonymous):

How simple. First add the two values on the left side and the 2 values inside the parentheses. Then just divide 48+54 by (8+9) and that will be you answer

OpenStudy (anonymous):

6

OpenStudy (anonymous):

@Juliannaaaaaa please don't give answers, as we are supposed to help them understand how to do them.

OpenStudy (anonymous):

plus it's against the open study guidelines, and the moderators will get really mad.

OpenStudy (anonymous):

6

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

but yes that is correct, it is 6

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

Mrs. Roberts plants tomatoes and radishes in the areas shown in the model below.

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Which expression is equivalent to the sum of the two areas? 5(7 + 4) 5(7 + 7) 7(5 + 4) 7(4 + 7) Mark this and return

OpenStudy (anonymous):

oops not mark this and return

OpenStudy (anonymous):

well the area is basically 35+28 so whichever of those equals that would be the answer

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

The third one

OpenStudy (anonymous):

63

OpenStudy (anonymous):

and oh

OpenStudy (anonymous):

Yes, that is 35 + 28. And like julianna said the only one that equals 63 is the 3rd but please, stop giving answers only.

OpenStudy (anonymous):

He's wrong u do not add those two

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Whaaaaaaa????????????

OpenStudy (anonymous):

Omg, yes you do. The area is 35+28

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Next question?

OpenStudy (anonymous):

Using the distributive property, an expression equivalent to 34 + 16 is 2(17 + 8). 17(2 + 8). 34 + (4)(4). (2)(17) + (4)(4).

OpenStudy (anonymous):

SO basically just first find the value of 34 + 16. That is 50. So which one of them equals 50? HInt, (4)(4) equals 16.

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

So based on that, the answer is the 3rd one.

OpenStudy (anonymous):

Hold on. This is distributive property. I did not see that.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

It would have to be A or B and once you solve it, you get A

OpenStudy (anonymous):

yeah this is distributive property

OpenStudy (anonymous):

all of the questions are about distributive property

OpenStudy (anonymous):

The other ones are right, lets move on

OpenStudy (anonymous):

Hello? @DeacontheVexHunter

OpenStudy (anonymous):

ok sorry i had to use the bathroom, i forgot to say brb

OpenStudy (anonymous):

Aida applied the distributive property to write the equivalent expressions below. 15 + 21 = 3(4 + 8) What is Aida’s error? Aida used 3 as the factor, which is not common to 15 and 21. Aida did not correctly divide by the common factor to get (4 + 8). Aida did not apply the correct operations to the expressions. Aida wrote two expressions that are not equivalent.

OpenStudy (anonymous):

Well here, they do equal each other, but i guess it means that she did not divide by the correct common factor, because if you look at every one of them you can eliminate everything but B

OpenStudy (anonymous):

b is the answer, thx

OpenStudy (anonymous):

Sari applied the distributive property using the greatest common factor to determine the expression that is equivalent to 84 + 40. Her work is shown below. Factors of 84: 1, 2, 3, 4, 6, 7, 12, 14, 21, 28, 42, 84 Factors of 40: 1, 2, 4, 5, 8, 10, 20, 40 84 + 40 = 2(42 + 20) What statement best describes Sari’s error? Sari did not use correct factors for 84 in the equation. Sari did not use correct factors for 40 in the equation. Sari did not use two equivalent expressions in the equation. Sari did not use the greatest common factor in the equation.

OpenStudy (anonymous):

okay, the answer is pretty obvious as there are bigger factors than 2. So that means?

OpenStudy (anonymous):

its not a or b??????????

OpenStudy (anonymous):

Yes, it is D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Do you understand the distributive property and GCF? no offense but for me this is some easy math.

OpenStudy (anonymous):

i suck at math, like i can bareley answer 2+2, lol, but im REALLY REALLY BAD AT MATH

OpenStudy (anonymous):

..... i have no comment.

OpenStudy (anonymous):

next question?

OpenStudy (anonymous):

i know im bad, LOL, but not really a little exageration

OpenStudy (anonymous):

Using the distributive property, an expression equivalent to 48 + 30 is 5(6 + 8). (6)(8) + (3)(10). (3)(16) + (6)(5). 6(8 + 5).

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

Sorry i was doing something. SO again it has to be either a or d, so which equals 48+30?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

YAY

OpenStudy (anonymous):

Jimmy applied the distributive property to write the equation below. 24 + 6 = 6(4 + 2) What is Jimmy’s error? Jimmy did not write the common factor in the correct place. Jimmy used 6 as the factor, which is not common to 24 and 6. Jimmy did not apply the correct operations to the expressions. Jimmy wrote two expressions that are not equivalent.

OpenStudy (anonymous):

you there???

OpenStudy (anonymous):

Agh this game ..... lol but it is the last one

OpenStudy (anonymous):

game? lol

OpenStudy (anonymous):

Which expression is shown using the model below?

OpenStudy (anonymous):

5(6 + 6) 5(6 + 5) 6(6 + 5) 6(6 + 10)

OpenStudy (anonymous):

HELP ONLY TEN MINUTES LEFT, PLEASE!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@triciaal please!!!!!!!!!!!!!!!!

OpenStudy (triciaal):

the common number is outside inside you have the light and the dark